Due to a lack of funding, the number of students enrolled at City College went from 6000 last year to 2000 this year. Find the percent decrease in enrollment. (Round to the nearest tenth of a percent, if necessary.)

Answer by Theo(13342)

enrollment went from 6000 to 2000.
the decrease in enrollment is 4000
4000/6000 = 4/6 = 2/3 = .6666667
the percent decrease is 66.67%
you start with 6000
multiply that by .6667 to get 4000.
subtract that from 6000 to get 2000.
6000 * (1-.66667) = .33333 * 6000 which is equal to 1999.98
if you use 2/3 instead of .66667, then you get the correct answer.
rounding to .66667 will leave some error due to rounding.
the percent decrease is .66667 * 1005 = 66.67%